calculate the angular momentum of an object rotating at 1 rad/s with a moment of intertia of 2 Kg m2.
Answer #1

The necessary formula to calculate the Angular Momentum (L) of a rotating object:

L = I \times \omega..........(i)

where

I = moment of inertia and \omega = angular velocity

From the given values in the question,

I = 2 kg.m2 and  \omega = 1 rad/sec.

Therefore, putting these values in eq(i), we get L = 2 * 1 = 2 kg.m2/sec.

Thus, the angular momentum of the rotating object is  2 kg.m^{2}/sec.